# AFIRE: Accurate and Fast Image Reconstruction Algorithm for Geometric-inconsistency Multispectral CT

### v1

@@ 4.2 Convergence analysis for the discrete setting
Consider the scheme 28 and 26, and denote the residual in each iteration by
@@
Under proper conditions, we have the following global convergence of algorithm 3.
@@
Assume that the conditions in lemma 2 hold and $\gamma<1$. The solution to 20 is unique.
Moreover, for any $k$, $\|\boldsymbol{\delta}^{k}\|_{2}\leq\zeta\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\|_{2}$, where $\zeta<(1-\gamma)/(1+\gamma)$. Then, the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 28 converges to the unique solution of 20.
@@
By mean value theorem and lemma 2, for any $\mathbf{f}$, we have the following estimate
@@
Since $\gamma<1$ and $\nabla\mathbf{K}(\hat{\mathbf{f}})$ has full column rank, let the $\mathbf{f}$ in section 4.2 be another solution if it exists, the uniqueness of the solution is immediately obtained by section 4.2.
@@
Hence, by 31 and 4.2, and the above inequality in order, we have
@@
$\leq\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}.$
@@
For the scheme 27, the residual $\boldsymbol{\delta}^{k}$ defined in 31 is vanishing. Following theorem 2, it is easy to obtain the convergence of algorithm 2.
@@
Assume that the conditions in lemma 2 hold, $\mathbf{P}$ is invertible and $\gamma<1$. Then the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 27 converges to the unique solution of 20.
@@
The proof of theorem 2 can be immediately used here, so we omit the detail. ∎
@@ 5 Numerical experiments
We carry out numerical experiments to validate the effectiveness and numerical convergence of the proposed algorithm AFIRE for solving the image reconstruction problem in the geometric-inconsistency DECT in the case of mildly full scan.
The compared algorithms include some existing methods such as the nonconvex primal-dual algorithm (NCPD) [8], the nonlinear Kaczmarz method using cyclic strategy (NKM) [14], and the iterative FBP reconstruction method (IFBP) [24]. Moreover, we consider a practical algorithm that uses linear interpolation firstly for the inconsistent data to obtain the data with consistency, and then employ the two-step DDD method to invert the obtained data, where the method for solving the nonlinear systems in the first step is chosen as Newton’s method. We refer to this algorithm as INTRPL for short.
@@ 5.1 Experiment settings
The tests are conducted on a workstation running Python equipped with an Intel i9-13900K @3.0GHz CPU and Nvidia RTX A4000 GPU. The experiment settings for the energy spectra and the MAC are presented as below. We simulate the discrete X-ray energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$ by an open-source X-ray spectrum simulator, SpectrumGUI [31]. Specifically, the used energy spectra pair, as displayed in fig. 1, contains an 80-kV spectrum and a filtered 140-kV spectrum of the GE maxiray 125 X-ray tube, where the filter is a copper of 1-mm width. The MAC of the basis materials, namely, water and bone, are sourced from [20] and used as the decomposition coefficients $\{\mathbf{b}_{d}\}_{d=1}^{D}$. Additionally, we employ two full-scan parallel-beam inconsistent geometries for low and high energy spectra, respectively.
@@
In algorithm 3 of AFIRE, as stated in remarks 3 and 4, we take $\hat{\mathbf{f}}=\boldsymbol{0}$, and implement the $\mathbf{P}_{q}^{+}$, $q=1,\ldots,Q$ by FBP method where the 1-bandwidth Ram–Lak filter is used. Moreover, INTRPL and IFBP also use the same FBP method, and NCPD is used to solve a total variation (TV)-regularized nonlinear least square problem with nonnegative constraint. For NCPD, the regularization parameter is selected as 00 and $10^{-6}$ for the noiseless and noisy data, respectively. Without loss of generality, the initial point for all tested algorithms are set to the zero vector.
@@ 5.2 Simulated noiseless data for the forbild phantom
In this test, a forbild phantom is composed of the truth water and bone basis images which are both discretized as $128\!\times\!128$ pixels on the square area $[-5,5]\!\times\![-5,5]$ $\text{cm}^{2}$ and take gray values over $[0,1]$, as shown in row 1 of fig. 5.
The X-ray paths under the different spectra displayed in fig. 1 are completely inconsistent. Specifically, for the 80-kV spectrum, we utilize 384 parallel projections uniformly sampled on $[-7.05,7.05]$ cm for each of the 384 views, where these views are uniformly distributed over the interval $[0,\pi)$ and can be seen as a mildly full scan. Under the filtered 140-kV spectrum, the scanning also has the same number of projections and views, but the views are uniformly distributed over the interval $[\pi/768,\pi/768+\pi)$.
@@
It is shown by the metrics of relative error that the proposed algorithm AFIRE has reached numerical convergence and reconstructed high-accuracy result in less than 50 iterations, thereby demonstrating its effectiveness and accuracy for inverting noiseless data. The results also confirm the convergence theory established in theorem 2. In contrast, NCPD and NKM are still far from convergence, and INTRPL can only produce low-accuracy result. Particularly, it is observed that IFBP does not converge.
@@ 5.3 Simulated noisy data for the realistic torso image
For this test, a real patient torso image is decomposed into the truth water and bone basis images which are both discretized as $256\!\times\!256$ pixels on the square area $[-5,5]\!\times\![-5,5]$ $\text{cm}^{2}$ and take gray values over $[0,1]$, as shown in row 1 of fig. 9.
Similarly, the X-ray paths under the different spectra shown in fig. 1 are also completely inconsistent. More precisely, under the 80-kV spectrum, 768 parallel-beam projections uniformly distributed on $[-7.05,7.05]$ cm are taken for each of 768 views, where these views are evenly sampled on the interval $[0,\pi)$. As to the filtered 140-kV spectrum, the same numbers of projections and views are utilized, but the views are uniformly distributed over the interval $[\pi/1536,\pi/1536+\pi)$. Moreover, the 34.3 dB noisy data is generated by adding Gaussian noise onto the noiseless data.
We plot the curves of metrics $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in fig. 6 within 100 iterations. It is obvious that the metrics of relative errors for the proposed algorithm AFIRE exhibit numerical convergence and relatively high accuracy in less than $15$ iterations, demonstrating its effectiveness and stability for addressing noisy data. On the contrary, NCPD and NKM have not converged yet; INTRPL can only produce low-accuracy result; IFBP is divergent after a couple of iterations.
@@
We also present the curves of metrics $\Delta_{\mathbf{f}}^{k}$ and $\Delta_{\mathbf{g}}^{k}$ in fig. 8, which definitely validate the numerical convergence of the proposed algorithm AFIRE.
@@ 6 Discussion
In this section, we discuss more details about the proposed algorithm AFIRE. Actually, this algorithm can be seen as a framework for solving the image reconstruction problem in MSCT, since there are two main adjustable modules, for instance, including the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$ in algorithm 3 of AFIRE. On the other hand, although this algorithm was originally designed to handle the reconstruction problem of geometric inconsistency between low and high X-ray energy spectra, it can be directly used to solve the problem in geometric-consistency situation. Hence, the flexibility and extensibility may enhance the applicability of the proposed algorithm.
To begin with, we introduce some basic experiment settings. The same energy spectra and MAC are applied as those described in section 5. Differently, the tested water and bone basis images are originated from the realistic CT images which are randomly selected from the dataset of LoDoPaB-CT [22], as shown in row 1 of figs. 11, 12 and 14. All the basis images are discretized as $362\!\times\!362$ pixels on the square area $[-15,15]\!\times\![-15,15]$ $\text{cm}^{2}$.
In the discussion of the two adjustable modules, under the 80-kV spectrum, we employ 1086 parallel projections evenly sampled on $[-21.15,21.15]$ cm for each of the 900 views, where these views are uniformly distributed over the interval $[0,\pi)$. Under the filtered 140-kV spectrum, the scanning uses the same numbers of projections and views, but the views are uniformly distributed over the interval $[\pi/1800,\pi/1800+\pi)$. With these configurations, we generate simulated noiseless data as before and apply algorithm 3 with selecting different modules to solve the corresponding reconstruction problems.
@@ 6.1 Selection of the special point $\hat{\mathbf{f}}$
Based on our insights in section 3, in the discrete (continuous) case, obviously, for the real constant set $\{C_{d}^{[q]}=0\}_{d,q=1}^{D,Q}$ ($\{C_{d}=0\}_{d=1}^{D}$), there exists $\hat{\mathbf{f}}=\boldsymbol{0}$ ($\hat{\boldsymbol{f}}=\boldsymbol{0}$) fulfilling the condition 23 (the condition 12). As we can imagine, those are almost the most simple examples satisfying the required conditions.
@@
As shown in fig. 10, we plot the curves of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in $100$ iterations.
@@ 6.2 Implementation of the approximated inverse $\mathbf{P}^{+}$
In section 5, we employed the FBP with 1-bandwidth Ram–Lak filter to implement the approximated inverse $\mathbf{P}^{+}$. Besides that, the proposed algorithm AFIRE also allows us to adopt different implementations for $\mathbf{P}^{+}$. Specifically, we further consider the other two different implementation methods: conjugate gradient method with 20 inner iterations (CG) and limited-memory BFGS method with 60 inner iterations (L-BFGS). Note that, the numbers of the inner iterations are not fixed, which can be modified on demand. All the implementations are running with Operator Discretization Library (ODL) [1].
@@
Moreover, as indicated in fig. 13, the metric curves of $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ also reveal the similar convergence behavior among those implementations.
@@
The implementation that requires inner iterations such as CG and L-BFGS achieves slightly lower error of $\text{RE}_{\mathbf{f}}^{k}$ than the used FBP. On the contrary, as evidenced by the comparison of computing time in table 1, FBP is faster than the other two implementations in single (outer) iteration and when the errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ reach $10^{-5}$. As proved in theorem 2, more precise implementation of $\mathbf{P}^{+}$ may improve convergence rate of the entire algorithm. However, it may increase the computational cost and reduce the efficiency. Therefore, there is a trade-off between the reconstructed accuracy and computational efficiency.
@@
  Single (outer) iteration $0.47$ 0.80 3.22
$\text{RE}_{\mathbf{f}}^{k}$ reaches $10^{-5}$ $13.99$ 20.70 86.81
$\text{RE}_{\mathbf{g}}^{k}$ reaches $10^{-5}$ $8.86$ 15.13 61.09
6.3 Extension to the geometric-consistency situation
From the derivation in section 3, we know that the proposed algorithm AFIRE can be naturally extended to the geometric-consistency case. To illustrate this, what is different from before is that the identical equidistant fan-beam geometry is used for both 80-kV and 140-kV spectra. There are totally 900 views uniformly distributed over $[0,2\pi)$ and 1086 equal-width detectors positioned on the opposing curved detector bank.
The proposed algorithm is used to reconstruct the basis images from the noiseless data. In the geometric-consistency case, the two-step DDD method can be used as a benchmark algorithm for comparison. In the first step, to estimate the basis sinograms, we use Newton’s method to solve the associated nonlinear systems in data domain, where the iteration number is 10 and the relative error achieves $10^{-14}$. In the second step, to reconstruct the basis images, the FBP with 1-bandwidth Ram–Lak filter is also applied to invert the obtained basis sinograms.
After running algorithm 3 for 100 iterations with setting $\hat{\mathbf{f}}=\boldsymbol{0}$, the obtained basis images and VMIs are shown in fig. 14, as well as the corresponding results gained by the two-step DDD method. Furthermore, the relative error curves are presented in fig. 15. These comparisons demonstrate that the performance of the proposed algorithm far exceeds the benchmark two-step DDD method. Hence, the proposed algorithm cannot only be successfully extended to deal with the geometric-consistency case of MSCT image reconstruction, but also achieve very high accuracy and efficiency.

### v2

@@ 4.2 Convergence analysis for the discrete setting
Considering the iterative scheme 29, and using 26, we denote a residual at each iteration by
@@
Under proper conditions, we have the following global convergence of the iterative scheme 29.
@@
Assume that the conditions in lemma 2 hold and $\gamma<1$. Then, the solution to 20 is unique.
Moreover, suppose that for any $k$, the residual in 33 satisfies
$\|\boldsymbol{\delta}^{k}\|_{2}\leq\zeta\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\|_{2},$ (34)
where the constant $\zeta<(1-\gamma)/(1+\gamma)$. Then, the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 29 converges to the unique solution of 20.
@@
By the mean value theorem and lemma 2, for any $\mathbf{f}$, we have the following estimate
@@
Letting $\bar{\mathbf{f}}^{*}$ be another solution if it exists, by section 4.2, we obtain
$\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\bar{\mathbf{f}}^{*}-\mathbf{f}^{*})\|_{2}\leq\gamma\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\bar{\mathbf{f}}^{*}-\mathbf{f}^{*})\|_{2}.$
Using the assumption of $\gamma<1$, it follows that
$\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\bar{\mathbf{f}}^{*}-\mathbf{f}^{*})\|_{2}=0.$
Since $\nabla\mathbf{K}(\hat{\mathbf{f}})$ has full column rank, we conclude that $\bar{\mathbf{f}}^{*}=\mathbf{f}^{*}$. Thus, the solution to 20 is unique.
@@
Hence, by 33, 34 and 4.2, and the above inequality in order, we have
@@
$\leq\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}.$ (36)
@@
Note that for the iterative scheme 27, the residual $\boldsymbol{\delta}^{k}$ in 33 is always vanishing. Following theorem 2, it is easy to obtain the convergence of the scheme 27, namely, algorithm 2. We present this result by the following corollary.
@@
Assume that the conditions in lemma 2 hold, $\mathbf{P}$ is invertible and $\gamma<1$. Then, the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 27 converges to the unique solution of 20.
@@
The proof of theorem 2 can be directly used here, so we omit the detail. ∎
Remark 9.
For the iterative scheme 29, the estimate of section 4.2 presents a linear convergence in the sense of $\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}$. Furthermore, by section 4.2, we have that
$\sigma_{\min}\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\|\mathbf{f}^{k+1}-\mathbf{f}^{*}\|_{2}\leq$ $\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k+1}-\mathbf{f}^{*})\|_{2}$
$\leq$ $\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}$
$\leq$ $\sigma_{\max}\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\mathbf{f}^{k}-\mathbf{f}^{*}\|_{2},$
which implies
$\|\mathbf{f}^{k+1}-\mathbf{f}^{*}\|_{2}\leq\kappa\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\mathbf{f}^{k}-\mathbf{f}^{*}\|_{2}.$ (37)
Here $\kappa(\cdot)=\sigma_{\max}(\cdot)/\sigma_{\min}(\cdot)$ denotes the condition number. Hence, if $\kappa\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\in(0,1)$, the Q-linear convergence of the proposed algorithm is established. Here the prefix “Q” stands for “quotient”.
Remark 10.
For the iterative scheme 29, a more precise implementation of $\mathbf{P}^{+}$ can generate a smaller residual $\boldsymbol{\delta}^{k}$ in 33. Then, there is a smaller $\zeta$ satisfying the condition of 34. From 37, it is easy to know that the smaller the $\zeta$, the faster the convergence rate of the algorithm.
@@ 5 Numerical experiments
We carry out numerical experiments to validate the effectiveness and numerical convergence of the proposed algorithm AFIRE for solving the image reconstruction problem in the geometric-inconsistent DECT in the case of mildly full scan.
The compared algorithms include some existing methods such as the nonconvex primal-dual (NCPD) algorithm [10], the nonlinear Kaczmarz method (NKM) using cyclic strategy [16], and the IFBP reconstruction method [26]. Moreover, we consider a practical algorithm that firstly processes the geometric-inconsistent data by using linear interpolation to obtain the geometric-consistent data with errors. Then the two-step DDD method is used to conduct the reconstruction. In the first step, the Newton’s method can be used to estimate the basis sinograms by solving a lot of very small-scale independent nonlinear systems along different X-ray paths. In the second step, to reconstruct the basis images, the FBP method is applied to invert the finally obtained basis sinograms. To show the intermediate results, the FBP method is also utilized to reconstruct the basis images from the estimated basis sinograms at each Newton iteration in the first step. We refer to this algorithm as INTRPL for short.
@@ 5.1 Experiment settings
The tests are conducted on a workstation running Python equipped with an Intel i9-13900K @3.0GHz CPU and Nvidia RTX A4000 GPU. The experiment settings for the energy spectra and the MAC are presented as below. We simulate the discrete X-ray energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$ by an open-source X-ray spectrum simulator, SpectrumGUI [34]. Specifically, the used energy spectra pair, as displayed in fig. 1, contains an 80-kV spectrum and a filtered 140-kV spectrum of the GE maxiray 125 X-ray tube, where the filter is a copper of 1-mm width. The MAC of the basis materials, namely, water and bone, are sourced from [22] and used as the decomposition coefficients $\{\mathbf{b}_{d}\}_{d=1}^{D}$. Notably, the use of two basis materials is a standard and physically well-justified practice in clinical DECT. Additionally, we employ two full-scan parallel-beam inconsistent geometries for low and high energy spectra, respectively.
@@
In algorithm 2 of AFIRE, as stated in remarks 6 and 8, we take $\hat{\mathbf{f}}=\boldsymbol{0}$, and implement the $\mathbf{P}_{q}^{+}$, $q=1,\ldots,Q$ by FBP method where the 1-bandwidth Ram–Lak filter is used. Moreover, INTRPL and IFBP also use the same FBP method, and NCPD is used to solve a total variation (TV)-regularized nonlinear least square problem with nonnegative constraint. For NCPD, the regularization parameter is selected as 0 and $10^{-6}$ for the noiseless and noisy data, respectively. Without loss of generality, the initial point for all tested algorithms are set to the zero vector.
@@ 5.2 Simulated noiseless data for the forbild phantom
In this test, a forbild phantom is composed of the truth water and bone basis images which are both discretized as $\!\times\!128$ 128 pixels on the square area $\!\times\![-5,5]$ [ - 5 , 5 ] $\text{cm}^{2}$ and take gray values over [0,1][0,1][ 0 , 1 ], as shown in row 1 of fig. 5.
The X-ray paths under the different spectra displayed in fig. 1 are completely inconsistent. Specifically, for the 80-kV spectrum, we utilize 384 parallel projections uniformly sampled on [7.05,7.05][-7.05,7.05][ - 7.05 , 7.05 ] cm for each of the 384 views, where these views are uniformly distributed over the interval $[0,\pi)$ and can be seen as a mildly full scan. Under the filtered 140-kV spectrum, the scanning also has the same number of projections and views, but the views are uniformly distributed over the interval $[\pi/768,\pi/768+\pi)$.
@@
It is shown by the metrics of relative error that the proposed algorithm AFIRE has reached numerical convergence and reconstructed high-accuracy result in less than 50 iterations, thereby demonstrating its effectiveness and accuracy for inverting noiseless data. The results also confirm the convergence theory established in theorem 2. In contrast, NCPD and NKM are still far from convergence, and INTRPL can only produce a low-accuracy result. Particularly, it is observed that IFBP does not converge. This is because IFBP, as a heuristic approach that designed for geometric-inconsistent case, lacks rigorous theoretical guarantees.
@@ 5.3 Simulated noisy data for the realistic torso image
For this test, a real patient torso image is decomposed into the truth water and bone basis images which are both discretized as $\!\times\!256$ 256 pixels on the square area $\!\times\![-5,5]$ [ - 5 , 5 ] $\text{cm}^{2}$ and take gray values over [0,1][0,1][ 0 , 1 ], as shown in row 1 of fig. 9.
Similarly, the X-ray paths under the different spectra shown in fig. 1 are also completely inconsistent. More precisely, under the 80-kV spectrum, 768 parallel-beam projections uniformly distributed on [7.05,7.05][-7.05,7.05][ - 7.05 , 7.05 ] cm are taken for each of 768 views, where these views are evenly sampled on the interval $[0,\pi)$. As to the filtered 140-kV spectrum, the same numbers of projections and views are utilized, but the views are uniformly distributed over the interval $[\pi/1536,\pi/1536+\pi)$. Moreover, the 34.3 dB noisy data is generated by adding Gaussian noise onto the noiseless data.
We plot the curves of metrics $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in fig. 6 within 100 iterations. It is obvious that the metrics of relative errors for the proposed algorithm AFIRE exhibit numerical convergence and relatively high accuracy in less than 151515 iterations, demonstrating its effectiveness and stability for addressing noisy data. On the contrary, NCPD and NKM have not converged yet; INTRPL can only produce a low-accuracy result; IFBP is divergent after a couple of iterations.
@@
We also present the curves of metrics $\Delta_{\mathbf{f}}^{k}$ and $\Delta_{\mathbf{g}}^{k}$ in fig. 8, which validate the numerical convergence of the proposed algorithm AFIRE.
@@ 6 Discussion
In this section, we discuss more details about the proposed algorithm AFIRE. Actually, this algorithm can be seen as a framework for solving the image reconstruction problem in MSCT, since there are two main adjustable modules, for instance, including the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$ in algorithm 2 of AFIRE. On the other hand, although this algorithm was originally designed to handle the reconstruction problem of geometric inconsistency between low and high X-ray energy spectra, it can be directly used to solve the problem in geometric-consistent situation. Hence, the flexibility and extensibility may enhance the applicability of the proposed algorithm.
To begin with, we introduce some basic experiment settings. The same energy spectra and MAC are applied as those described in section 5. Differently, the tested water and bone basis images are originated from the realistic CT images which are randomly selected from the dataset of LoDoPaB-CT [24], as shown in row 1 of figs. 11, 12 and 14. All the basis images are discretized as $\!\times\!362$ 362 pixels on the square area $\!\times\![-15,15]$ [ - 15 , 15 ] $\text{cm}^{2}$.
In the discussion of the two adjustable modules, under the 80-kV spectrum, we employ 1086 parallel projections evenly sampled on [21.15,21.15][-21.15,21.15][ - 21.15 , 21.15 ] cm for each of the 900 views, where these views are uniformly distributed over the interval $[0,\pi)$. Under the filtered 140-kV spectrum, the scanning uses the same numbers of projections and views, but the views are uniformly distributed over the interval $[\pi/1800,\pi/1800+\pi)$. With these configurations, we generate simulated noiseless data as before and apply algorithm 2 with selecting different modules to solve the corresponding reconstruction problems.
@@ 6.1 Selection of the special point $\hat{\mathbf{f}}$
Based on our insights in section 3, in the discrete (continuous) case, obviously, for the real constant set $\{C_{d}^{[q]}=0\}_{d,q=1}^{D,Q}$ ($\{C_{d}=0\}_{d=1}^{D}$), there exists $\hat{\mathbf{f}}=\boldsymbol{0}$ ($\hat{\boldsymbol{f}}=\boldsymbol{0}$) fulfilling the condition 23 (the condition 12). As we can imagine, those are almost the simplest examples that satisfying the required conditions.
@@
As shown in fig. 10, we plot the curves of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in 100100100 iterations.
@@ 6.2 Implementation of the approximated inverse $\mathbf{P}^{+}$
In section 5, we employed the FBP with 1-bandwidth Ram–Lak filter to implement the approximated inverse $\mathbf{P}^{+}$. Besides that, the proposed algorithm AFIRE also allows us to adopt different implementations for $\mathbf{P}^{+}$. Specifically, we further consider the other two different implementation methods: conjugate gradient (CG) method with 20 inner iterations and limited-memory BFGS (L-BFGS) method with 60 inner iterations. Note that, the numbers of the inner iterations are not fixed, which can be modified on demand. All the implementations are running with Operator Discretization Library (ODL) [1].
@@
Moreover, as indicated in fig. 13, the metric curves of $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ reveal the similar convergence behavior among those implementations. The implementation that requires inner iterations, such as CG and L-BFGS, achieves a slightly lower $\text{RE}_{\mathbf{f}}^{k}$ than FBP after the same number of outer iterations. In contrast, as evidenced by the comparison of running time in table 1, the FBP implementation is faster in a single (outer) iteration and when the errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ reach $10^{-5}$.
As described in remark 10, a more precise implementation of $\mathbf{P}^{+}$ may improve the convergence rate. However, it may increase the computational cost or reduce the efficiency. Therefore, there is a trade-off between the reconstructed accuracy and computational efficiency.
@@
  Single (outer) iteration 0.470.470.47 0.80 3.22
$\text{RE}_{\mathbf{f}}^{k}$ reaches $10^{-5}$ 13.9913.9913.99 20.70 86.81
$\text{RE}_{\mathbf{g}}^{k}$ reaches $10^{-5}$ 8.868.868.86 15.13 61.09
6.3 Extension to the geometric-consistent situation
From the derivation in section 3, we know that the proposed algorithm AFIRE can be naturally extended to the geometric-consistent case. To illustrate this, what is different from before is that the identical equidistant fan-beam geometry is used for both 80-kV and 140-kV spectra. There are totally 900 views uniformly distributed over $[0,2\pi)$ and 1086 equal-width detectors positioned on the opposing curved detector bank.
The proposed algorithm is used to reconstruct the basis images from the noiseless data. In the geometric-consistent case, the two-step DDD method can be used as a benchmark algorithm for comparison. In the first step, to estimate the basis sinograms in data domain, we use Newton’s method to solve a lot of very small-scale independent nonlinear systems along different X-ray paths. When the iteration number is 101010, the relative error has arrived at $10^{-14}$. In the second step, to reconstruct the basis images, the FBP with 1-bandwidth Ram–Lak filter is applied to invert the finally obtained basis sinograms. To show the intermediate results, the FBP method is also utilized to reconstruct the basis images from the estimated basis sinograms at each Newton iteration in the first step.
After running algorithm 2 for 100 iterations with setting $\hat{\mathbf{f}}=\boldsymbol{0}$, the obtained basis images and VMIs are shown in fig. 14, as well as the corresponding results gained by the two-step DDD method. Furthermore, the relative error curves are presented in fig. 15. These comparisons demonstrate that the performance of the proposed algorithm far exceeds the benchmark two-step DDD method. Hence, the proposed algorithm cannot only be successfully extended to deal with the geometric-consistent case of MSCT image reconstruction, but also achieve very high accuracy and efficiency.